### v1

# RetrySQL: text-to-SQL training with retry data for self-correcting query generation

Abstract.
The text-to-SQL task is an active challenge in Natural Language Processing. Many existing solutions focus on using black-box language models extended with specialized components within customized end-to-end text-to-SQL pipelines. While these solutions use both closed-source proprietary language models and coding-oriented open-source models, there is a lack of research regarding SQL-specific generative models. At the same time, recent advancements in self-correcting generation strategies show promise for improving the capabilities of existing architectures. The application of these concepts to the text-to-SQL task remains unexplored. In this paper, we introduce RetrySQL, a new approach to training text-to-SQL generation models. We prepare reasoning steps for reference SQL queries and then corrupt them to create retry data that contains both incorrect and corrected steps, divided with a special token. We continuously pre-train an open-source coding model with this data and demonstrate that retry steps yield an improvement of up to 4 percentage points in both overall and challenging execution accuracy metrics, compared to pre-training without retry data. Additionally, we confirm that supervised fine-tuning with LoRA is ineffective for learning from retry data and that full-parameter pre-training is a necessary requirement for that task. We showcase that the self-correcting behavior is learned by the model and the increase in downstream accuracy metrics is a result of this additional skill. Finally, we incorporate RetrySQL-trained models into the full text-to-SQL pipeline and showcase that they are competitive in terms of execution accuracy with proprietary models that contain orders of magnitude more parameters. RetrySQL demonstrates that self-correction can be learned in the text-to-SQL task and provides a novel way of improving generation accuracy for SQL-oriented language models.
text-to-SQL, SQL, retry data, self-correction, continued pre-training, large language model
ccs: Computing methodologies Natural language processing
1. Introduction
The task of translating natural language questions to SQL queries is a major challenge for machine learning models. The complexity stems from the need of relating often ambiguous user input to abstract entities, relations and values that are present in relational databases (Li et al., 2023). Even prominent Large Language Models (LLMs), such as GPT-4o (Team, 2024b) or Gemini 1.5 (Team, 2024a), struggle with approaching human performance in leading text-to-SQL benchmarks: BIRD (Li et al., 2023) and SPIDER 2.0 (Lei et al., 2024). The same is true for models tuned specifically for coding tasks (Li et al., 2024; Chen et al., 2024; Talaei et al., 2024). Thus, there exists a need for more advanced solutions that can bridge that gap and provide reliable SQL queries even in difficult real-world scenarios.
The text-to-SQL task can be divided into three main steps (Maamari et al., 2024): retrieval, generation and correction. Many existing approaches try to tackle these steps at the same time, in a single end-to-end pipeline (Maamari et al., 2024; Talaei et al., 2024; Pourreza and Rafiei, 2023, 2024). In this work, we focus only on the generation step and show how it can be improved with a novel approach to model pre-training.
Figure 1. RetrySQL overview. (a) Reasoning step generation. For each SQL query in the training dataset, we generate a series of reasoning steps using GPT-4o. (b) Preparation of retry data. For each set of reasoning steps, we apply random perturbations, treated as errors, by replacing some steps with different ones. We follow these errors with special [BACK] tokens and amend them with correct steps. (c) We take an open-source LLM and continue its pre-training with training examples that contain retry data injected into reasoning steps. The resulting RetrySQL-trained model learns the ability to self-correct, which improves its capabilities in generating correct SQL queries from natural language questions.
Specifically, we teach the model to self-correct during the generation itself. While previous work did use self-correction in the sense of post-processing, applied in the correction step (Pourreza and Rafiei, 2023), we enforce the self-correcting behavior at an earlier point. This sort of active knowledge-based self-correction is an ongoing research area when it comes to LLMs (Ye et al., 2024a, b). While recent work in slow thinking reasoning systems, such as DeepSeek-R1 (Team, 2025), shows that self-correction can be learned in a reinforcement learning setup, other lines of research suggest that is is possible to obtain the self-correction ability with specific data augmentations and a standard auto-regressive pre-training objective (Ye et al., 2024a, b). It has been shown that augmenting training data for grade-school math solution generation with so-called retry data leads to increased generation accuracy (Ye et al., 2024b). The applicability of this approach to other tasks and models has not been explored as of yet.
We introduce RetrySQL, a novel text-to-SQL generation module training paradigm that incorporates retry data in the training process and teaches the resulting model to self-correct. RetrySQL first augments the training data with reasoning steps that explain the sequence of operations required for obtaining the solution SQL query (Fig. 1a). Then, retry data is generated by corrupting the order of these reasoning steps (Fig. 1b). The retry data is incorporated into the training examples and we perform continued pre-training of an open-source coding-oriented LLM, which results in a RetrySQL-trained model that is capable of self-correction (Fig. 1c). We present a set of experiments across multiple strategies of generating reasoning corruptions. We demonstrate that using retry data yields superior generation results when compared to training data with error-free reasoning steps.
To demonstrate that the self-correcting behavior is indeed learned by the model, we provide an analysis of output token confidence. We show that the max softmax score is on average lower for tokens before correction than for those after. Similarly, incorrect tokens display higher variance of softmax scores across beam search passes than the corrected ones. This illustrates that the model becomes uncertain as it makes a mistake and then self-corrects itself with higher confidence.
The standard approach to adapting open-source LLMs to specific tasks is supervised fine-tuning (SFT) with Low-Rank Adaptation (LoRA) (Hu et al., 2022), which is a byproduct of large model sizes and limited compute resources for most machine learning practitioners. We show that SFT with LoRA is insufficient for learning the self-correction ability from retry data and that full- parameter continued pre-training is a necessary requirement for that task.
Our results corroborate the recent findings regarding the self-correction ability in language models (Ye et al., 2024b), demonstrating that the improvement in generation accuracy coming from the inclusion of retry data in pre-training is a universal law. It is applicable not only to the grade school math reasoning problem and GPT-2, but also to the text-to-SQL domain and larger, more modern Transformer-based decoder-only models. These findings suggest that retry data could be adapted to even more domains, especially if reasoning steps can be added to the training examples.
While the main focus of our work is on the SQL generation step in isolation, we also showcase that relatively small, 1.5B-parameter open-source coding models trained with our RetrySQL paradigm are competitive with much larger closed-source proprietary LLMs when used as a part of the full text-to-SQL pipeline. We share all code for retry data generation, as well as model training and evaluation [^1] https://github.com/allegro/RetrySQL.
In summary, our key contributions are the following:
  • We introduce RetrySQL, a novel text-to-SQL training paradigm that makes use of reasoning steps enhanced with retry data.
  • We show that using retry data in pre-training is beneficial to the generation process, as indicated by the Execution Accuracy metric calculated for the BIRD benchmark dataset.
  • We demonstrate that RetrySQL-trained models have the ability to self-correct as they generate reasoning steps for the output SQL queries.
  • We show that full-parameter continued pre-training is necessary for retry data in the text-to-SQL task, as opposed to the typical LoRA-based supervised fine-tuning.
  • We illustrate that within a simple end-to-end text-to-SQL pipeline, RetrySQL-trained 1.5B-parameter open-source coding models are competitive with proprietary models such as GPT-4o-mini and GPT-4o.
2. Related work
Early text-to-SQL methods relied on sequence-to-sequence frameworks, using models like Graph Neural Networks, Recurrent Neural Networks, and pre-trained Transformers for encoding queries and schemas (Cai et al., 2021; Cao et al., 2021; Hwang et al., 2019), while employing slot-filling or auto-regressive decoding to generate SQL queries (Choi et al., 2020; Wang et al., 2020). Recently, the field has shifted with the emergence of LLMs, which are currently leading in the most popular benchmarks (Shi et al., 2024). While initial efforts were focused on optimizing prompt designs that leveraged in-context learning (Nan et al., 2023; Gao et al., 2023) and multi-stage prompting (Pourreza and Rafiei, 2023), the current state-of-the-art is represented by LLM-based pipelines. These latest approaches integrate LLMs in more complex sequences of processing stages, with separate components for schema linking, self-correction, self-debugging, and self-consistency (Talaei et al., 2024; Pourreza et al., 2024; Lee et al., 2024; Maamari et al., 2024; Sun et al., 2024).
Compared to closed-source model prompting approaches, open-source model fine-tuning for the text-to-SQL task remains relatively unexplored (Shi et al., 2024). Many of the existing works favor parameter-efficient fine-tuning (PEFT) over full-parameter fine-tuning due to the former’s superior training efficiency and lower training costs (Shi et al., 2024; Chen et al., 2024; Zhang et al., 2024a, b). While the majority of practitioners choose to use powerful general-purpose LLMs as their base models (Pourreza and Rafiei, 2024; Xie et al., 2024; Shi et al., 2024), promising results have also been shown by adapting coding LLMs to the text-to-SQL domain (Li et al., 2024; Chen et al., 2024; Talaei et al., 2024), demonstrating that starting from a model already heavily pre-trained on coding tasks, with SQL-related training data, leads to higher performance in benchmark evaluations.
The bulk of text-to-SQL research aims to utilize LLMs as a tool integrated in an end-to-end pipeline, and only a very small effort has been dedicated to studying the capabilities of such models in the context of SQL generation alone. It has been shown that language models trained to follow chain of thought (CoT) steps excel at solving problems that involve math and symbolic reasoning (Mirzadeh et al., 2024; Sprague et al., 2024; Ye et al., 2024a, b), but these findings have not been validated as of yet in the context of the text-to-SQL task. Recent work in the domain of LLM theory paves the way for the discovery of LLM universal laws, pertaining to learning language structures (Allen-Zhu and Li, 2024a), knowledge storage and manipulation (Allen-Zhu and Li, 2024b, c, d), learning from mistakes and the ability to self-correct (Ye et al., 2024a, b). We aim to validate the applicability of the latter to coding LLMs and the text-to-SQL downstream task.
3. Methodology
In this section, we describe our RetrySQL training paradigm in detail. We augment the BIRD benchmark dataset (Li et al., 2023) (Section 3.1) with synthetically generated reasoning steps (Section 3.2). Then, we define the retry data generation process, in which reasoning steps are corrupted with random errors and then corrected (Section 3.3).
3.1. Training data
In order to acquire a sizable training dataset for the text-to-SQL task, we utilized the existing training data from the BIRD benchmark (Li et al., 2023). The dataset includes 9428 examples, each consisting of the database name, natural language question, external evidence, and the ground truth SQL query (SQLite dialect). In addition, the metadata for each database is available as well, consisting of a full list of tables, columns and table relations. We discovered that relations for one table, mondial_geo, are defined incorrectly. We excluded it from our pipeline, which left us with 9135 training examples.
For the generation process, we needed to incorporate the schema information together with the question and external knowledge. To this end, we parsed the ground truth queries and prepared the matching Data Definition Language (DDL) statements, which served as the schema linking data. Importantly, unless stated otherwise, we incorporated so-called perfect schema linking in our experiments (i.e. no redundant links). We were interested primarily in studying the SQL generation process in isolation, without the additional task of finding relevant schema connections. We matched column and table names in each ground truth SQL query to the corresponding database metadata and built minimal required schema links.
We used DDL for schema representation because it provides a concise notation that includes table and column names, together with data types and relations. Moreover, it keeps an SQL-focused context for the model, without needing to explain specific data formats in the prompt. This approach is commonly used in existing text-to-SQL pipelines (Shi et al., 2024).
3.2. Reasoning step generation
Previous work showed that the usage of retry data in model pre-training is effective only if we also instruct the model to follow a chain of reasoning steps (Ye et al., 2024b). To this end, we needed to procure reasoning steps for each of our training examples. The BIRD training dataset does not contain this data, so we used GPT-4o for generating synthetic reasoning steps (Fig. 1a). Enhancing language model training data with synthetically generated components is a newly emerging trend (Abdin et al., 2024). We used a prompt that highlighted the need of reasoning steps being in a format resembling solution reasoning chains from the dataset used in previous research on self-correction (Ye et al., 2024a) (Fig. S3). We verified the correctness of the output formatting and also the semantic validity for a small subset of all examples, consisting of 100 instances sampled uniformly at random, thus representing a wide range of databases and query difficulties. We found that there were no cases with erroneous reasoning steps. Consequently, we assumed that the full dataset is similarly error-free. Full manual verification was not necessary, since we did not aim for the reasoning steps to be perfectly accurate. We ultimately wanted to generate SQL queries, and the reasoning steps were meant to serve as an additional training signal to the verified ground truth from the BIRD dataset.
3.3. Retry data generation
We generated SQL-specific retry data by corrupting the solution steps prepared beforehand (following (Ye et al., 2024b)) (Fig. 1b). We considered several variants of perturbations: forward single (denoted as FS), forward and back single (FBS), forward multiple (FM), forward and back multiple (FBM). Given a sequence of reasoning steps of length $N$, for each step $r_{i}$ in that sequence we select uniformly at random (with probability $p_{retry}$) another step $r_{error}\in S$, where $S$ is a set of candidate corruptions. The selection is done either once (for FS and FBS) or multiple times (for FM and FBM). For FS and FM, $S$ consists of elements $r_{i+1},...\ ,r_{N}$ (i.e. future steps). For FBS and FBM, $S$ contains elements $r_{1},...\ ,r_{i-1},r_{i+1},...\ ,r_{N}$ (i.e. future and past steps). After selecting an element $r_{error}$ from $S$, we follow it with the [BACK] token and then with the correct $r_{i}$ itself. As such, each $r_{i}$ can be replaced with the following:
  • ($r_{error}$, [BACK], $r_{i}$), for FS and FBS,
  • ($r_{error}$, [BACK], … , $r_{error}$, [BACK], $r_{i}$), for FM and FBM.
We generated training dataset variants for FS, FBS, FM, FBM and $p_{retry}\in\{0.1,0.2,0.3,0.4,0.5\}$. To steer the training, we introduced additional tokens [CONTEXT], [QUESTION], [REASONING] and [SQL]. We combined the database schema, external knowledge, reasoning steps and the ground truth SQL query using these tokens as delimiters (Fig. S2).
4. Experiments
In this section, we outline the experimental setup for our study. We delineate the preliminary linear probing task (Section 4.2), and then describe the baseline models that we used in our experiments (Section 4.1). We present details regarding the experiment for SFT with LoRA (Section 4.3). We explain our inference procedure, as well as evaluation metrics used to measure the effectiveness of all models (Section 4.4).
4.1. Baseline models
We evaluated several baselines in addition to the models trained with retry data. To measure the gap between our solution and proprietary models, we evaluated zero-shot performance of GPT-4o-mini [^1] API version: 2023-03-15-preview, GPT-4o11footnotemark: 1, Gemini-1.5-flash [^2] Stable version: 002 and Gemini-1.5-pro22footnotemark: 2. For these models, we used the prompt template and inference configuration that followed the BIRD baselines (Li et al., 2023), including zero-shot CoT and $temperature=0.0$. We set ${max\_output\_tokens=2048}$ to avoid truncated outputs.
For our experiments with retry data, we chose the open-source OpenCoder 1.5B (Huang et al., 2024) model. Such a small model size allowed us to more effectively utilize our compute resources and sped up the experimentation process. Our goal was to first and foremost validate the effectiveness of the RetrySQL training paradigm, for which a larger number of parameters was not necessary. The OpenCoder model was initially trained with coding corpora that already contained SQL data. As such, when we further continued training this model, the SQL-specific knowledge didn’t have to be learned from scratch and the model could focus strictly on the specifics of text-to-SQL generation with reasoning steps.
For the details on the training setup and hyperparameters, see Appendix A.1.
4.2. Linear probing dataset
Before conducting the pre-training experiments with retry data, we first validated if the OpenCoder model has an innate, hidden capability of distinguishing correct and incorrect reasoning steps. If that were the case, then providing retry data at training time would allow the model to unlock the ability to self-correct. It has been shown previously that models pre-trained on grade school math data with correct solution steps exhibit regretful patterns in their internal states (Ye et al., 2024b).
In order to verify the above hypothesis, we designed a preliminary linear probing task. We parsed the retry data and categorized the training examples based on the presence of the [BACK] token. Each reasoning step $r{{}_{i}}$ was marked as either:
  • incorrect, if it was followed by the [BACK] token;
  • correct, if it was followed by another step.
Then, we took the original BIRD data samples and extended them with reasoning step sequences ending with either correct or incorrect steps, again with the addition of special [CONTEXT], [QUESTION] and [REASONING] tokens to divide the input sections. We used the Retry FS 0.3 dataset variant as the source for the reasoning steps. In this way, we extracted $15k$ examples in total, keeping the proportion between correct and incorrect instances balanced.
We used this data to train a classification model, in which a binary classification head categorized correct and incorrect reasoning steps (for more details, see Appendix A.4).
Table 1. Execution Accuracy for models evaluated in a zero-shot scenario. Models were prompted using the BIRD baseline prompts, as described in 4.1 and assuming perfect schema linking.
Model name EXsimple EXmoderate EXchallenging EXoverall
GPT-4o 72.37 ± 0.51 53.16 ± 1.31 43.59 ± 1.79 63.73 ± 0.15
GPT-4o-mini 47.31 ± 0.4 24.31 ± 1.12 21.38 ± 1.38 37.91 ± 0.25
Gemini-1.5-pro 75.59 ± 0.93 59.87 ± 0.67 59.03 ± 2.32 69.27 ± 0.94
Gemini-1.5-flash 74.68 ± 1.04 59.01 ± 0.18 56.14 ± 1.87 68.20 ± 0.84
OpenCoder 1.5B 40.04 ± 0.20 16.90 ± 0.48 7.45 ± 0.91 29.96 ± 0.07
4.3. Supervised fine-tuning with LoRA
In addition to continued pre-training, we also examined the scenario of SFT with retry data, reflecting a practical situation where an open-source pre-trained model is adapted to enhance reasoning capabilities or to address a specific task. Our objective was to evaluate whether fine-tuning with retry data is as effective as continued pre-training. We focused on PEFT, specifically LoRA (Hu et al., 2022), which is a technique commonly used in practice. LoRA fine-tunes a limited set of trainable parameters, keeping the original pre-trained weights frozen. Typically, for Transformer-based dense models, LoRA is applied either to just the query/value matrices or to all linear layers of the network (Hu et al., 2022; Mao et al., 2024). Since in the fine-tuning process we added new special tokens (namely [REASONING] and [BACK]), both input and output embeddings needed to be trained as well. We fine-tuned the base model with retry data directly, we did not start with a model pre-trained using only error-free data (contrary to (Ye et al., 2024b)). In the latter configuration, LoRA is incapable of leveraging the retry data without destroying the previously learned knowledge.
4.4. Inference process and evaluation metrics
For the purpose of measuring the effectiveness of retry data in text-to-SQL generation, we utilized the Execution Accuracy (EX) metric introduced in the BIRD benchmark. In all experiments we used the development dataset provided by BIRD for evaluation. It contains a total of 1534 examples. The data format is the same as the training set, with the addition of a difficulty value (simple, moderate, challenging) for each example. During inference, we used the same format as during training, but omitted the part of each sequence after the [REASONING] token. Thus, we wanted the model to first generate the reasoning steps, and then the [SQL] token, followed by the actual SQL query.
During inference we used the best checkpoint of each trained model variant (except for SFT with LoRA, where we used the last checkpoint, following  (Ye et al., 2024b)). We used beam search multinomial sampling with 4 beams as a decoding strategy (following (Ye et al., 2024b)), with $temperature=0.5$, $top\_k=50$ and $top\_p=1.0$. We limited the number of new tokens to 1024. Each evaluation example was processed 5 times, for the purpose of measuring the model variance. The results were post-processed by removing the [SQL] and preceding tokens, leaving only the SQL query.
We then calculated the EX metric in the following manner. First, the generated SQL query as well as the ground truth query for a given question were executed in an SQLite database containing the development data. Then, resulting sets of rows were compared and the ratio of matching rows was saved. Finally, the match ratios for all examples were averaged. We report four EX values: EXoverall, EXsimple, EXmoderate, EXchallenging. They correspond to an overall EX over all examples, or over just the simple, moderate or challenging ones, respectively.
5. Results
In this section, we assess the effectiveness of RetrySQL for training SQL generation models. We showcase the baseline results for both proprietary and open-source LLMs (Section 5.1). We demonstrate through linear probing that the baseline OpenCoder model can recognize incorrect reasoning steps (Section 5.2). We then show that training with retry data improves Execution Accuracy compared to training with error-free reasoning steps (Section 5.3). We verify that SFT with LoRA is incapable of training models that utilize retry data (Section 5.4). Finally, we explain the self-correction behavior with an analysis of model confidence around the [BACK] tokens (Section 5.5).
5.1. Zero-shot baselines
We evaluated OpenCoder 1.5B , as well as several proprietary LLMs, in zero-shot mode (Tab. 1). We observe that OpenCoder 1.5B falls behind state-of-the-art models, with an EXoverall score gap ranging from 7.95 percentage points (p.p.) for GPT-4o-mini, to 39.31 p.p. for Gemini-1.5-pro. The same is true for the detailed metrics as well, which overall signifies relatively poor performance in the text-to-SQL task. These results indicate that, in the zero-shot setting, even a recent open-source coding-oriented model such as OpenCoder is not able to reach the performance of general-purpose proprietary solutions.
5.2. Detecting regretful patterns through linear probing
Figure 2. t-SNE projection of OpenCoder’s internal state embeddings for the linear probing task. Blue points represent embeddings corresponding to correct reasoning steps, while orange points indicate embeddings for incorrect steps. The clusters of orange points indicate that the OpenCoder model differentiates a large portion of the incorrect steps from the correct ones, highlighting the innate, yet hidden, ability to detect mistakes in the reasoning process.
We took the baseline OpenCoder model and performed a linear probing experiment with its frozen weights. The linear probing model detected incorrect steps with average $balanced\_accuracy=82\%$ and $f1\_score=71\%$. Since the detection accuracy was significantly higher than $50\%$ (random guess), we can conclude that the probing signals most likely came from the pre-trained weights, and not from the fine-tuned classification layer (Ye et al., 2024a). To support our findings, we visualize the internal state embeddings of OpenCoder using a t-SNE projection (Fig. 2). The plot illustrates the separability of the model’s internal states when it comes to predicting correct versus incorrect reasoning steps. These results demonstrate that OpenCoder has an innate capability to self-correct, which is in line with previous research positing that this ability is a universal law for all Transformer models (Ye et al., 2024b). This justifies the usage of retry data, which should enable the model to backtrack as it generates an incorrect step and then retry once again.
Table 2. Execution Accuracy for the continued pre-training case with RetrySQL. All results are expressed in percentages, with mean and standard deviation over 5 multinomial beam search generations. The best results are marked in bold. Results with retry data that improve upon the error-free training are indicated with an underline.
Dataset variant EXsimple EXmoderate EXchallenging EXoverall
– (zero-shot) 40.04 ± 0.20 16.90 ± 0.48 7.45 ± 0.91 29.96 ± 0.07
error-free 62.70 ± 0.07 43.53 ± 0.14 39.45 ± 0.28 54.71 ± 0.08
Retry FS 0.1 65.84 ± 0.00 44.09 ± 0.11 36.83 ± 0.34 56.52 ± 0.04
Retry FS 0.2 68.22 ± 0.12 45.47 ± 0.14 40.28 ± 0.34 58.70 ± 0.09
Retry FS 0.3 68.00 ± 0.00 44.91 ± 0.26 43.31 ± 0.28 58.68 ± 0.06
Retry FS 0.4 66.57 ± 0.11 44.22 ± 0.32 34.62 ± 0.28 56.79 ± 0.11
Retry FS 0.5 66.98 ± 0.18 44.96 ± 0.29 37.79 ± 0.28 57.56 ± 0.12
5.3. Retry data improves SQL generation metrics
To test if the OpenCoder model can learn the ability to self-correct, we continued the pre-training process with reasoning-enhanced BIRD training data, both error-free and with retry data. Compared to the zero-shot OpenCoder 1.5B baseline, the model continuously pre-trained with error-free data yielded an impressive improvement in generation accuracy metrics (Tab. 2): $\sim$23 p.p. for simple examples, $\sim$27 p.p. for moderate ones, and 32 p.p. for challenging instances, which resulted in an overall increase of $\sim$25 p.p.. These results are expected, as during training we provided the model with previously unseen domain-specific text-to-SQL training samples.
The retry data results show us that models continuously pre-trained with such corrupted samples lead to improved accuracy metrics when compared to the error-free continued pre-training (Tab. 2, Tab. S1). Out of four approaches to retry data preparation, the FS variant proved to be the best overall. Thus, here we present findings only for the FS variant. For the results pertaining to the other variants (i.e. FM, FBS, FBM), see Appendix A.2. The highest improvement in overall generation accuracy, 3.99 p.p., was observed for $p_{retry}$=0.2. The second best increase, 3.97 p.p., resulted from employing the Retry FS 0.3 dataset.
While the EXoverall metric does show the effectiveness of using retry data for improving text-to-SQL generation accuracy in general, it does not show the full picture. Looking at the EXsimple score, we see that the biggest improvement stems from using the dataset with $p_{retry}$=0.2 (increase of 5.52 p.p.). For the moderate cases, the biggest improvement of the EX score equals 2.15 p.p. (once again for $p_{retry}$=0.2). The biggest increase in the EXchallenging metric, 4.00 p.p., resulted from pre-training with the Retry FS 0.3 dataset. Previous research on self-correcting generation with retry data postulated that the advantage of the self-correction ability can be observed especially for complex out-of-distribution evaluation examples, which required the longest solution reasoning sequences (Ye et al., 2024b). Our results show that for the text-to-SQL task, the improvements are more even across difficulty levels. This is a consequence of how the BIRD dataset is constructed and of our approach to generating reasoning steps (see Section 3.2) - the number of operations required to explain SQL queries is not perfectly correlated with the difficulty level. There is a significant overlap in the complexity of SQL queries across difficulty levels in the BIRD development dataset (Fig. S1). As such, many challenging examples can be solved with relatively short queries, which do not necessitate the model to generate long reasoning step sequences. Examples that do require more reasoning are simply spread out across all difficulty levels, which explains the relatively even increases of Execution Accuracy observed in our experiments. These results show that the self-correction ability introduced by pre-training with retry data is an effective way of improving the accuracy of SQL generation.
Table 3. Execution Accuracy for SFT with LoRA. We performed LoRA fine-tuning using only the Retry FS 0.3 variant. All results are expressed in percentages, with mean and standard deviation over 5 multinomial beam search generations.
LoRA rank EXsimple EXmoderate EXchallenging EXoverall
– (error-free) 62.70 ± 0.07 43.53 ± 0.14 39.45 ± 0.28 54.71 ± 0.08
${r=8}$ 14.98 ± 0.20 7.07 ± 0.32 5.24 ± 0.34 11.67 ± 0.17
${r=16}$ 56.22 ± 0.27 30.34 ± 0.21 23.72 ± 0.70 45.32 ± 0.17
${r=32}$ 56.76 ± 0.22 30.60 ± 0.49 22.62 ± 0.28 45.62 ± 0.15
${r=64}$ 60.02 ± 0.16 36.38 ± 0.16 30.34 ± 0.00 50.07 ± 0.09
${r=128}$ 58.90 ± 0.16 36.98 ± 0.17 23.03 ± 0.34 48.88 ± 0.09
${r=256}$ 58.96 ± 0.09 38.45 ± 0.17 28.55 ± 0.34 49.88 ± 0.12
5.4. SFT with LoRA is ineffective for training with retry data
For the experiments that evaluated SFT with LoRA, we used the Retry FS 0.3 dataset as the training data. While it scored marginally lower than the Retry FS 0.2 variant in the overall EX metric, the high performance for the challenging examples marked it as a preferable dataset.
It is evident that SFT with LoRA is not effective when it comes to learning the self-correction ability from retry data (Tab. 3). None of the tested LoRA ranks resulted in models that improved upon the error-free training. For rank 8 the model was not able to learn almost anything from the training data ($EX_{overall}$ equal to 11.67%). For higher ranks some knowledge was transferred (overall accuracy up to 50.07%), but the limited number of trainable parameters was too low to effectively preserve both the reasoning step generation and self-correction abilities at the same time.
Consequently, to utilize the retry data and teach a coding LLM to self-correct in the text-to-SQL task, it is necessary to perform full-parameter continued pre-training. This finding supports observations made in previous work regarding the capability of LoRA-trained models to learn from retry data (Ye et al., 2024b).
While the usage of LoRA is the only way for most practitioners to fine-tune sizable (70B parameters and more) LLMs, the fact that a small 1.5B-parameter model trained with RetrySQL achieves very good results indicates that continued pre-training of smaller models with the addition of reasoning-based retry data might be a viable alternative.
Figure 3. Distribution of token confidence before and after [BACK] tokens. (a) Mean of max token confidence across 10 beam search passes. It can be seen that the confidence score is on average much higher for tokens after the [BACK] token, indicating that the model is uncertain as it makes mistakes, but is confident after self-correction. (b) Standard deviation of max token confidence across 10 beam search passes. The variance of model predictions is much higher as it makes mistakes than after self-correction. Both of these results show that the self-correcting behavior is indeed learned by a model trained with RetrySQL. We used RetrySQL-FS-0.3 for obtaining these results.
5.5. Does RetrySQL know that it makes mistakes?
The EX results show us that training with RetrySQL increases the number of correct SQL queries compared to models trained without retry data (Tab. 2). However, it could be the case that the additional tokens present in reasoning steps in retry data simply act as robust augmentation and the model does not learn to self-correct. The underlying model behavior needs to be studied in more detail. To this end, we analyzed the confidence scores returned by the model trained with the Retry FS 0.3 dataset (denoted further as RetrySQL-FS-0.3) in proximity (radius of 10 tokens) to [BACK] tokens. We took max softmax scores from these tokens, and then calculated the mean and standard deviation per token across 10 multinomial beam search passes. Finally, we averaged these metrics separately for tokens before and after each [BACK] token.
It is evident that the mean of max confidence scores for predicted tokens differs between tokens preceding the [BACK] token and those after it (Fig. 3a). In other words, as the model is making mistakes, it is less confident in its predictions than after it self-corrects itself and starts to generate correct tokens. This shows that the ability to self-correct is an active part of the generation process.
Similarly, the standard deviation of confidence scores across beam search passes differs between tokens before and those after the [BACK] token (Fig. 3b). For the incorrect tokens, the variance is on average much higher than for the ones after the self-correction. This indicates a reduction in model uncertainty - before the [BACK] token each beam search pass returns significantly different results, the model is not decided what to choose. Conversely, after self-correction, the results become much more consistent and certain - the model catches the error and commits to the correction.
Both of these results clearly show that the self-correcting ability is a learned behavior, resulting from the inclusion of retry data in the training process. For examples of model output that contains reasoning steps with self-correction, see Appendix A.7.
Table 4. Execution Accuracy for the full end-to-end text-to-SQL pipeline. All results are expressed in percentages, with mean and standard deviation over 5 multinomial beam search generations.
Model name EXsimple EXmoderate EXchallenging EXoverall
GPT-4o 61.62 ± 3.83 42.74 ± 2.34 40.48 ± 1.78 54.99 ± 0.32
GPT-4o-mini 42.12 ± 1.06 18.06 ± 0.47 17.65 ± 0.38 32.53 ± 0.72
Gemini-1.5-pro 66.88 ± 0.23 48.15 ± 0.62 51.86 ± 0.58 59.79 ± 0.31
Gemini-1.5-flash 64.69 ± 0.12 45.86 ± 0.24 49.52 ± 0.57 57.56 ± 0.08
RetrySQL-FS-0.2 59.81 ± 0.04 37.46 ± 0.09 35.17 ± 0.00 50.72 ± 0.00
RetrySQL-FS-0.3 60.28 ± 0.05 38.36 ± 0.14 36.00 ± 0.28 51.36 ± 0.05
5.6. RetrySQL-trained model in an end-to-end text-to-SQL pipeline
All experiments presented thus far focused strictly on the generation step, with perfect pre-computed schema linking. In that setting, we found that the best RetrySQL-trained models, denoted further as RetrySQL-FS-0.2 and RetrySQL-FS-0.3 (trained with Retry FS 0.2 and Retry FS 0.3, respectively), achieved EXoverall scores of 58.70% and 58.68% (Tab. 2). These metrics are significantly higher than the result for the proprietary GPT-4o-mini model (by $\sim$20.8 p.p.), and quite close to the score for the GPT-4o-model (short of it by $\sim$5 p.p.) (Tab. 1). However, in a real text-to-SQL pipeline, perfect schema linking is not available. In order to validate if RetrySQL-FS models are competitive with existing proprietary models in the full end-to-end pipeline setting, we performed an additional set of experiments. For the complete description of the pipeline, together with details on our schema linking methodology, see Appendix A.3.
In the full pipeline setting, we observed that the RetrySQL-trained models remain competitive with much larger models (Tab. 4). The EXoverall scores for RetrySQL-FS-0.2 and RetrySQL-FS-0.3 are still much higher than the result for GPT-4o-mini (by $\sim$18-19 p.p.). At the same time, the gap to GPT-4o is smaller in the full pipeline setting: 3.63 p.p. (for RetrySQL-FS-0.3), compared to $\sim$5 p.p. for the perfect schema linking setting.
Crucially, it must be highlighted that the RetrySQL-FS models are relatively small (1.5B parameters). The aforementioned proprietary GPT-4o-mini and GPT-4o LLMs are estimated to be much larger ($\sim$8B and $\sim$200B parameters (Abacha et al., 2025), respectively).
These results indicate that using retry data in conjunction with our RetrySQL method produces 1.5B-parameter models that are competitive with much larger proprietary models, as measured by the BIRD EX metric. This is a promising outcome, showing that incorporating self-correction in the generation stage might be a way forward for future text-to-SQL end-to-end pipelines.
6. Limitations
In our experiments we used the training data from the BIRD benchmark, which contains a limited number of training examples. This is different than what has been done in previous work on self-correction with retry data (Ye et al., 2024b). There, the training examples were generated on demand as the training went on, to fill a preset number of training steps. We did not have a setup for generating synthetic data in that way, and had to rely on the curated training examples from BIRD. This might explain the discrepancy in the relative effectiveness of retry data - with the on-demand synthetic examples, the observed improvements in generation accuracy were in the range of 10-16 p.p. for the hardest examples. However, it is important to keep in mind that a direct comparison to our results is not obvious, as the problem setting of grade school math is very different to our text-to-SQL task. In addition, the metric used in that work was a direct measure of correctness, while in our case we used an indirect Execution Accuracy metric computed in relation to database values, which were not present in the training data. Furthermore, as indicated previously (see Section 5.3), the difficulty of examples in the BIRD development set does not correlate with reasoning length. However, despite these differences, the effectiveness of our approach is still evident, as the EX metric was noticeably improved thanks to the addition of retry data in the training examples. We leave synthetic training data generation as a topic for future work.
For the full pipeline experiments, we utilized only a relatively simple LLM-based schema linking stage and did not include a correction stage at the end. This is not an ideal strategy, as there are many optimizations that could be applied to these stages. However, the main part of our research focused on the generation stage and these other elements remained out of scope for us. Moreover, because the RetrySQL training paradigm teaches the generation model to self-correct, the additional correction step becomes less important. We leave building a fully optimized end-to-end text-to-SQL pipeline, with the most recent approaches to schema linking and query selection, as a topic for future research.
7. Conclusions
In this paper, we presented RetrySQL, a novel approach to training text-to-SQL generation models. Our solution utilizes reasoning steps with retry data in the training examples, which teaches the generation model to self-correct itself as it produces its output. We show that using such data for the continued pre-training of a coding LLM leads to improved Execution Accuracy metrics when compared to models pre-trained without retry data (increase of $\sim$4 p.p. both overall and for the challenging examples). In addition, we confirm previous observations related to using SFT with LoRA for the purpose of training with retry data. We provide an explainability analysis for our results - we show that as the RetrySQL model makes mistakes, it is less confident in its predictions than after it self-corrects itself. Finally, we showcase that incorporating RetrySQL-trained 1.5B-parameter models into a relatively simple end-to-end text-to-SQL pipeline produces results that are competitive with much larger closed-source proprietary LLMs such as GPT-4o-mini and GPT-4o. We hope that our RetrySQL training paradigm will lead to further developments in text-to-SQL models, especially in the context of self-correcting generation.
Appendix A Appendix
A.1. Training details
All models were trained with NVIDIA A100 80GB, utilizing 2 GPUs. The effective batch size equaled 128, due to gradient accumulation being set to 4 and the per-device batch size set to 16. We used dynamic per-batch right padding, with sequence length being padded to a multiple of 8. We utilized the AdamW optimizer with the following hyperparameters (inspired by (Li et al., 2024)): $\beta_{1}$=0.9, $\beta_{2}$=0.95, $epsilon$=1e-08, $weight\_decay$=0.1. In all experiments, we trained for 5 epochs, with $learning\_rate$=5.0e-05 and a cosine learning rate schedule. For optimizing the memory usage, we employed the DeepSpeed framework (Rasley et al., 2020) with stage Zero-2.
For experiments regarding SFT with LoRA, we utilized the same hardware and hyperparameters as described above. Following previous work (Ye et al., 2024b), we examined this set of low-rank configurations: ${r\in\{8,16,32,64,128,256\}}$. We applied $lora\_dropout=0.01$ and set $\alpha=2r$, following the common best practices (Hu et al., 2022).
A.2. Full RetrySQL results
In this section, we provide a full evaluation of all retry data variants (Tab. S1), together with a visualization for the distribution of SQL query complexity across data instances in the BIRD development dataset (Fig. S1).
The full results demonstrate that the FS retry data variant is the most effective out of the four that were considered in our study. While the EXoverall metric for the other variants improves upon the baseline error-free training in most cases, none of the results match the findings for the FS variant. This is even more evident for the detailed difficulty metrics, for which the FM, FBS and FBM variants are either worse than the FS variant, or worse than the error-free baseline altogether (see especially the EXchallenging metric). These results showcase that in the text-to-SQL task, retry data in reasoning steps needs to be sampled in the forward direction once per step, since other strategies are to a large extent not as efficient.
Table S1. Execution Accuracy for the continuous pre-training case with RetrySQL, with all retry data variants. The OpenCoder 1.5B model was used as the starting point for all trainings. All results are expressed in percentages. Since we found that the model variance across multinomial beam search passes is relatively low for the FS datasets (Tab. 2), we did not calculate standard deviations for the remaining variants. The best results are marked in bold. Results with retry data that improve upon the error-free training are indicated with an underline.
Dataset variant EXsimple EXmoderate EXchallenging EXoverall
– (zero-shot) 40.04 16.90 7.45 29.96
error-free 62.70 43.53 39.45 54.71
Retry FS 0.1 65.84 44.09 36.83 56.52
Retry FS 0.2 68.22 45.47 40.28 58.70
Retry FS 0.3 68.00 44.91 43.31 58.68
Retry FS 0.4 66.57 44.22 34.62 56.79
Retry FS 0.5 66.98 44.96 37.79 57.56
Retry FM 0.1 63.68 43.97 35.86 55.08
Retry FM 0.2 64.97 41.38 39.31 55.41
Retry FM 0.3 66.81 44.18 37.24 57.17
Retry FM 0.4 64.76 41.81 36.55 55.15
Retry FM 0.5 57.51 37.28 28.28 48.63
Retry FBS 0.1 66.70 43.53 34.48 56.65
Retry FBS 0.2 66.59 44.40 34.48 56.84
Retry FBS 0.3 67.03 45.26 35.86 57.50
Retry FBS 0.4 68.32 43.10 35.86 57.63
Retry FBS 0.5 66.16 42.67 35.17 56.13
Retry FBM 0.1 66.38 41.81 37.24 56.19
Retry FBM 0.2 65.95 43.97 33.79 56.26
Retry FBM 0.3 67.46 43.53 37.93 57.43
Retry FBM 0.4 66.05 44.18 31.72 56.19
Retry FBM 0.5 64.65 40.09 30.34 53.98
Figure S1. Distribution of SQL query complexity in the BIRD development dataset. There is a significant overlap in the number of level-1 expressions in the SQL syntax tree across difficulty levels defined in the BIRD development dataset. Due to our reasoning generation strategy (see Section 3.2), the number of these expressions is a proxy for the number of reasoning steps. We parsed the ground truth SQL queries with the SQLGlot Python library (SQLGlot, [n. d.]) and extracted level-1 elements from the corresponding syntax trees.
A.3. Text-to-SQL pipeline
In this section we describe our full text-to-SQL pipeline and provide more insights into all ablations and experiments conducted during the design process for the best schema linking approach.
A.3.1. Pipeline description
Our full text-to-SQL pipeline consisted of two modules:
  • schema linking: an LLM-based schema linker was executed in order to connect the correct tables and columns with the natural language question, thus retrieving the essential context for the generation step;
  • generation: RetrySQL-FS-0.2 and RetrySQL-FS-0.3 models were used to generate a sequence of reasoning steps and the final SQL query.
A.3.2. Schema linking
Schema linking is one of the most critical steps within the full text-to-SQL pipeline. While recent advancements suggest that it might be omitted in the case of using LLMs for the query generation (Maamari et al., 2024), feeding the model with too much data can provoke hallucinations and increase the inference cost due to a large number of processed input tokens (Chung et al., 2025). Moreover, the full database schema might simply not fit into the context window of the model.
Inspired by leading solutions in the BIRD benchmark (Pourreza and Rafiei, 2023, 2024), we treated schema linking as a separate task and designed a bespoke experimentation framework solely for the purpose of evaluating selected schema linking solutions. We extracted ground truth schema links from the BIRD development set and then investigated several schema linking algorithms. The tested algorithms can be grouped in the following categories:
  • heuristic based methods: algorithms that do not use any machine learning (ML) techniques, and instead try to link table and column names either by exact matching or by employing edit distance thresholds;
  • embedding based methods: ML-based algorithms that embed column and table representations and try to match them in the vector space via embedding similarity;
  • LLM-based methods: LLM-based algorithms that aim to find the correct schema linking by prompting an LLM to solve this specific task.
We compared these approaches by measuring the following metrics:
  • FP - false positive rate, indicates the proportion of irrelevant columns retrieved over the total number of columns;
  • recallcol - fraction of correctly retrieved columns per example, averaged over all test examples;
  • recalllink - proportion of test examples for which all required columns for perfect schema linking were retrieved.
Both embedding-based and LLM-based methods were evaluated in zero-shot mode (no training or fine-tuning involved), using the BIRD development set. For embedding-based models, we performed k-NN search with k=30. We prepared the queries by joining the natural language question and external knowledge, while documents were represented and stored as the set of all possible combinations of table name - column name - column description - column data format in the target database.
Our results show that LLM-based schema linking approaches represent the current best method for solving the schema linking problem (Tab. S2). Among the tested cloud-based LLMs, Gemini-1.5-pro achieved the best results: 0.37 FP, 0.92 recallcol, 0.77 recalllink. Thus, Gemini-1.5-pro was chosen for the schema linking stage of our pipeline. Empirically, we observed that specifying foreign and primary key metadata in the database schema boosted the results for all implemented methods. For completeness, we provide the prompt used for retrieving the schema linking and an example of the model’s output (Fig. S4).
Table S2. Results for the schema linking experiments. The best performing approach, as measured by all metrics, was the LLM-based schema linking with Gemini-1.5-pro. For FP, lower values are better. For recallcol and recalllink higher values are better. The best results are marked in bold.
Method FP recallcol recalllink
Heuristic-based methods
exact-matching 0.75 0.54 0.16
edit-distance=1 0.79 0.62 0.25
edit-distance=2 0.84 0.66 0.27
edit-distance=3 0.9 0.72 0.33
Embedding-based methods
OpenAI - text-embeddings-ada-002 0.73 0.64 0.28
OpenAI - text-embeddings-3-large 0.88 0.81 0.54
Snowflake - arctic-embed-m (Merrick, 2024) 0.75 0.6 0.23
Snowflake - arctic-embed-l (Merrick, 2024) 0.89 0.78 0.48
LLM-based methods
GPT-4o 0.38 0.92 0.76
GPT-4o-mini 0.43 0.85 0.56
Gemini-1.5-pro 0.37 0.92 0.77
Gemini-1.5-flash 0.39 0.91 0.72
A.4. Linear probing experiments
In this section we provide more details regarding our linear probing experiments, which were inspired by a probing task originally introduced in previous work (Ye et al., 2024a) (specifically, the $can\_next(A)$ task).
A.4.1. Experiment setup
We prepared the linear probing model by taking the OpenCoder-1.5B weights continuously pre-trained with error-free reasoning steps and replacing the existing head with a linear one. The new head mapped the $2044$-dimensional vector of the last token in the input sequence to a single sigmoid-activated neuron for the binary classification task. Unlike (Ye et al., 2024a), we did not introduce any small rank-r update on the input (embedding) layer. We trained this classification model using a machine equipped with 2 x NVIDIA A100 80GB, with effective $batch\_size$=128 and $learning\_rate$=1e-4. We used the AdamW optimizer with $\beta_{1}$=0.9, $\beta_{2}$=0.95 and utilized early stopping.
We trained the model in a 5-fold cross-validation setup, by retaining $80\%$ of the dataset for training and using the rest for validation. We used the best checkpoints to compute $balanced\_accuracy$ and $f1\_score$ (reported for the incorrect class).
A.4.2. Discussion
By performing the linear probing experiment we demonstrated that a coding-oriented LLM, previously pre-trained with only error-free reasoning steps, exhibits regretful patterns during inference for input examples that contain an incorrect sequence of reasoning steps. It is possible to leverage the probing results by applying the classifier described above to guide the SQL generation process. After generating each solution sentence, the next probing could be used to determine whether the model knows that it has made a mistake and, if so, it could be reverted to the end of the previous sentence and regenerate from that point. Previous work showed that this method can increase generation accuracy in mathematical reasoning tasks, at the cost of increased inference complexity (Ye et al., 2024b). We did not perform such an analysis, since we treated the probing experiment as a preliminary sanity check for using retry data in the text-to-SQL task.
A.5. Training data example
[CONTEXT] CREATE TABLE ’games’
 ’games_year’ INTEGER DEFAULT null
– External Knowledge: games refer to id; during the 90s refers to games_year BETWEEN 1990 AND 1999; – Using valid SQLite and understanding External Knowledge, answer the following questions for the tables provided above. [QUESTION] How many Olympics games were held during the 90s? [REASONING] Define the main table in the FROM clause: FROM games. Define the filtering condition in the WHERE clause: WHERE games_year BETWEEN 1990 AND 1999. Select the column to be included in the final result: SELECT COUNT(games_year). [SQL] SELECT COUNT(games_year) FROM games WHERE games_year BETWEEN ’1990’ AND ’1999’;
Figure S2. Example of a training data sample used in our experiments. It consists of the following elements: DDL statements for schema representation, external knowledge and question extracted from the BIRD metadata, reasoning steps generated as described in Section 3.2 and the ground truth SQL query. These components are separated by special tokens to guide the model in the learning process.
A.6. Prompts
You are a SQL expert. When I provide you with a SQL query, your task is to describe step-by-step how a person would create such a query. Follow the standard SQL execution order. Write the steps from the perspective of someone constructing the query. If the query includes subqueries, describe them step-by-step in the same detailed manner as the main query before referencing them. Each step should represent a distinct operation. Make the operations as granular as possible.
Here is the standard SQL execution order you should follow for your explanation. If a given clause is not present in the query, skip it without mentioning its absence:
1. FROM clause (including JOINs).
2. WHERE clause.
3. GROUP BY clause.
4. HAVING clause.
5. SELECT clause.
6. ORDER BY clause.
7. LIMIT clause.
For each query I provide:
1. Explain the query step by step in plain language.
2. Ensure that each step corresponds to one small, logical operation.
3. Use clear and concise language for each operation.
4. Each step should be provided in a single line (use single newline character between steps).
Here is an example query for your reference:
SELECT T1.name, T1.email, SUM(T3.amount) AS total_sales FROM Customers AS T1 INNER JOIN Orders AS T2 ON T1.customer_id = T2.customer_id LEFT JOIN OrderDetails AS T3 ON T2.order_id = T3.order_id WHERE T2.order_date >= ’2023-01-01’ AND T2.order_date <= (SELECT order_date FROM Orders WHERE order_id = ’1’ ORDER BY order_date DESC LIMIT 1) GROUP BY T1.name, T1.email ORDER BY total_sales DESC
The expected step-by-step breakdown for the above query:
Define the main table in the FROM clause: FROM Customers AS T1.
Define the first JOIN operation: INNER JOIN.
Define the table to join: Orders AS T2.
Define the join condition: ON T1.customer_id = T2.customer_id.
Define the second JOIN operation: LEFT JOIN.
Define the table to join: OrderDetails AS T3.
Define the join condition: ON T2.order_id = T3.order_id.
Define the main filtering condition in the WHERE clause: WHERE T2.order_date >= ’2023-01-01’.
Add the additional filtering condition in the WHERE clause : AND T2.order_date <= (subquery).
Define the main table in the subquery’s FROM clause: FROM Orders.
Define the main filtering condition in the subquery’s WHERE clause: WHERE order_id = ’1’.
Select the column to be included in the subquery result: SELECT order_date.
Order the subquery results by the specified column: ORDER BY order_date DESC.
Limit the subquery results: LIMIT 1.
Complete the filtering condition in the WHERE clause: AND T2.order_date <= (SELECT order_date FROM Orders WHERE order_id = ’1’ ORDER BY order_date DESC LIMIT 1).
Group the results by the specified columns: GROUP BY T1.name, T1.email.
Select the columns to be included in the final result: SELECT T1.name, T1.email, SUM(T3.amount) AS total_sales.
Order the results by the specified column: ORDER BY total_sales DESC.
Now, I will provide you with a query, and I expect you to respond in this format:
${\{sql\_query\}}$
Figure S3. Prompt used for generating reasoning steps.
– Database Schema: CREATE TABLE ‘yearmonth‘
 ‘customerid‘ INTEGER not null,
 ‘date‘ TEXT not null,
 ‘consumption‘ REAL null,
 FOREIGN KEY (‘customerid‘) REFERENCES ‘customers‘ (‘customerid‘),
 FOREIGN KEY (‘customerid‘) REFERENCES ‘customers‘ (‘customerid‘),
 PRIMARY KEY (‘date‘, ‘customerid‘)
CREATE TABLE ‘customers‘
 ‘customerid‘ INTEGER not null PRIMARY KEY,
 ‘currency‘ TEXT null
– External Knowledge: Pays in euro = Currency = ’EUR’
– Based on Database Schema provided above and understanding External Knowledge, your task is to select table-column pairs (called schema links) most relevant to the given Question. – Question: Among the customers who paid in euro, how many of them have a monthly consumption of over 1000? Choose the relevant table-column pairs after thinking step by step:
 "schema_links": [
   "table_name": "customers",
   "columns": [ "customerid", "currency" ]
   "table_name": "yearmonth",
   "columns": [ "consumption", "customerid", "date" ]
Figure S4. Example of a prompt used for the schema linking step in our proposed text-to-SQL full pipeline. The prompt is marked with yellow background, while the response is denoted with blue background. Database schemas are represented in form of DDL statements as described in Section 3.1. We used the LangChain (LangChain, [n. d.]) library to query LLMs, leveraging its Python dataclass serialization support. Note: the injected instructions for the output JSON format are not included in this example.
A.7. RetrySQL generation examples
[CONTEXT]
CREATE TABLE ‘cards‘
 ‘id‘ INTEGER not null PRIMARY KEY,
 ‘artist‘ TEXT null,
 ‘cardkingdomfoilid‘ TEXT null,
 ‘cardkingdomid‘ TEXT null
– External Knowledge: designed by Aaron Miller refers to artist = ’Aaron Miller’; are icredibily powerful refers to cardKingdomFoilId is not null AND cardKingdomId is not null
– Using valid SQLite and understanding External Knowledge, answer the following questions for the tables provided above.
[QUESTION]
Of all the cards that are designed by Aaron Miller, how many of them are incredibly powerful?
[REASONING]
Define the main table in the FROM clause: ‘FROM cards‘. Add the additional filtering condition in the WHERE clause: ‘AND cardKingdomFoilId IS NOT NULL‘. [BACK]
Define the main filtering condition in the WHERE clause: ‘WHERE artist = ’Aaron Miller’‘.
Add the additional filtering condition in the WHERE clause: ‘AND cardKingdomFoilId IS NOT NULL‘. Select the column to be included in the final result: ‘SELECT COUNT(id)‘. [SQL]
SELECT COUNT(id) FROM cards WHERE artist = ’Aaron Miller’ AND cardKingdomFoilId IS NOT NULL AND cardKingdomId IS NOT NULL"
Figure S5. RetrySQL generation example with prompt and completion. The prompt is marked with yellow background, while the response is denoted with blue background. The model corrects itself after erroneously generating a reasoning step in its CoT process. Specifically, the model generates an additional filtering condition when the main one has not been declared yet. After generating the [BACK] token, it retries and correctly generates the main filtering condition, and then proceeds until the end of the reasoning chain, finally producing a correct SQL query.
[CONTEXT]
CREATE TABLE ‘cards‘
 ‘id‘ INTEGER not null PRIMARY KEY,
 ‘artist‘ TEXT null,
 ‘convertedmanacost‘ REAL null,
 ‘side‘ TEXT null
– External Knowledge: doesn’t have multiple faces refers to side IS NULL; illustrator refers to artist
– Using valid SQLite and understanding External Knowledge, answer the following questions for the tables provided above.
[QUESTION]
Among the cards that doesn’t have multiple faces on the same card, who is the illustrator of the card art that has the highest cost of converted mana?
[REASONING]
Define the main table in the FROM clause: ‘FROM cards‘. Define the main filtering condition in the WHERE clause: ‘WHERE side IS NULL‘. Select the column to be included in the final result: ‘SELECT artist‘. Limit the results: ‘LIMIT 1‘. [BACK] Add the additional filtering condition in the WHERE clause: ‘AND convertedManaCost = (subquery)‘. Define the main table in the subquery’s FROM clause: ‘FROM cards‘. Complete the filtering condition in the WHERE clause: ‘AND convertedManaCost = (SELECT MAX(convertedManaCost) FROM cards)‘. Select the column to be included in the final result: ‘SELECT artist‘. Limit the results: ‘LIMIT 1‘. [BACK] Order the results by the specified column: ‘ORDER BY artist DESC‘. Limit the results: ‘LIMIT 1‘. [SQL]
SELECT artist FROM cards WHERE side IS NULL AND convertedManaCost = ( SELECT MAX(convertedManaCost) FROM cards ) ORDER BY artist DESC LIMIT 1
Figure S6. RetrySQL generation example with prompt and completion. The prompt is marked with yellow background, while the response is denoted with blue background. The model corrects itself after erroneously trying to halt the reasoning chain too early. This happens two times along the process. The retry mechanism allows the model to understand the mistakes and correct itself. Specifically, the model corrects itself by adding additional WHERE and ORDER operations. After each correction, the model proceeds to generate the rest of the reasoning chain, finally producing a correct SQL query.
[CONTEXT]
CREATE TABLE ‘yearmonth‘
 ‘customerid‘ INTEGER not null,
 ‘date‘ TEXT not null,
 ‘consumption‘ REAL null,
 FOREIGN KEY (‘customerid‘) REFERENCES ‘customers‘ (‘customerid‘),
 FOREIGN KEY (‘customerid‘) REFERENCES ‘customers‘ (‘customerid‘),
 PRIMARY KEY (‘date‘, ‘customerid‘)
CREATE TABLE ‘customers‘
 ‘customerid‘ INTEGER not null PRIMARY KEY,
 ‘currency‘ TEXT null
– External Knowledge: Pays in euro = Currency = ’EUR’.
– Using valid SQLite and understanding External Knowledge, answer the following questions for the tables provided above. [QUESTION]
Among the customers who paid in euro, how many of them have a monthly consumption of over 1000?
[REASONING]
Define the main table in the FROM clause: ‘FROM customers AS T1‘. Define the JOIN operation: ‘INNER JOIN‘. Define the table to join: ‘yearmonth AS T2‘. Select the column to be included in the final result: ‘SELECT COUNT(T1.CustomerID)‘. [BACK]
Define the join condition: ‘ON T1.CustomerID = T2.CustomerID‘.
Select the column to be included in the final result: ‘SELECT COUNT(T1.CustomerID)‘. [BACK]
Define the main filtering condition in the WHERE clause: ‘WHERE T2.Consumption > 1000‘.
Select the column to be included in the final result: ‘SELECT COUNT(T1.CustomerID)‘. [BACK]
Add the additional filtering condition in the WHERE clause: ‘AND T1.Currency = ’EUR’‘.
Select the column to be included in the final result: ‘SELECT COUNT(T1.CustomerID)‘. [SQL]
SELECT COUNT(T1.CustomerID) FROM customers AS T1 INNER JOIN yearmonth AS T2 ON T1.CustomerID = T2.CustomerID WHERE T2.Consumption > 1000 AND T1.Currency = ’EUR’
Figure S7. RetrySQL generation example with prompt and completion. The prompt is marked with yellow background, while the response is denoted with blue background. The model corrects itself after erroneously trying to halt the reasoning chain too early. This happens three times along the process. The retry mechanism allows the model to understand the mistakes and correct itself. In particular, the model corrects itself by adding additional JOIN and WHERE operations. After each correction, the model proceeds to generate the rest of the reasoning chain, finally producing a correct SQL query.